# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v16

@@ 9.13 Ideal Coverage
Adding noise to the input can suppress unspecified input errors and improve the portion of specified versus unspecified input uncertainties. Table 2 shows the measured minimal required noise to achieve ideal coverage for FFT transformations at FFT order $18$ for different signals and sine functions.
Figure 57 illustrates the error deviations and uncertainty means when applying an input noise of $10^{-3}$ to a Linear signal when using the Library sine function result in ideal coverage for all the transformations:
  • As expected, the result uncertainty means for the forward transformations increase with FFT order $L$ as $\sqrt{2}^{L}$.
  • As expected, the result uncertainty means for the reverse transformations decrease with FFT order $L$ as $1/\sqrt{2}^{L}$.
  • As expected, the result uncertainty means for the roundtrip transformations remains equal to the corresponding input uncertainties of $10^{-3}$.
  • As expected, the result error deviations for the forward and reverse transformations remain constant at $1$.
Figure 58 presents the corresponding histogram. As expected, the normalized errors for the forward and reverse transformations follow Normal distributions, while those for the roundtrip transformation are Delta distributed around zero, indicating perfect recovery of the input uncertainties. According to Figure 58, adding either Gaussian or white noise to input results in the same Gaussian distribution of the normalized errors.
Additionally, the result uncertainty means for both forward and reverse transformations are linearly proportional to the input uncertainties, as expected since FFT is a linear algorithm [12].
The range of ideal coverage depends on how accurately the specified input uncertainty represents the actual input noise. For Linear signals computed using the Library sine function, Figure 59 and 60 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively. Ideal coverage corresponds to the region where the error deviation equals $1$. Outside this region, proper coverage cannot be achieved for the reverse transformation. Because uncertainties grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region. Furthermore, as numerical errors increase with computational load, the range of input noise that produces ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders, visually beyond FFT order $25$ for the reverse transformation, ideal coverage may no longer be achievable. Although FFT is widely regarded as one of the most robust numerical algorithms, and generally insensitive to input errors, it can still fail due to numerical errors in the Library sine function. Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
In contrast, for Linear signals computed using the Quart sine functions, Figure 61 and 62 show the results for the forward and reverse transformations, respectively. The ideal coverage region is significantly larger, and proper coverage is also achieved in adjacent regions. The forward transformation exhibits a broader ideal coverage region than the reverse transformations. While the forward transformation shows consistent over-estimation of the value errors in the proper coverage region, the reverse transformation demonstrates both over-estimation and under-estimation of the value errors in the proper coverage region. Figure 62 shows that the rounding errors balances the over-estimation of value errors in the Quart sine function in Figure 38, so that the error deviations within the proper coverage region differ from $1$ only marginally.
Adding noises has another obvious benefit: Different implementations of the variance arithmetic have almost identical result when the input noise is larger than $10^{-16}$. For example, Figure 61 and 62 are almost identical using either C++, or Java, or Python.
9.14 Input Uncertainty Recovery
With input uncertainty unspecified, the value error waveform for the roundtrip transformation is qualitatively similar to that of the the reverse transformation, with error deviations close to $1$, such as in Figure 41, 43, 47, and 53, even when proper coverage can no longer achieve for both forward and reverse transformations in Figure 55. With input uncertainty unspecified, roundtrip transformation can always recover input uncertainty statistically.
With input uncertainty specified, roundtrip transformation can always recover input uncertainty numerically, and the quality of the recover increases with the quality of input uncertainty specification. For example, when using the Library sine function:
  • Figure 57 shows that the error deviations for the roundtrip transformation is less than $1$, suggesting that each input uncertainty is recovered numerically. These error deviations increases with FFT orders, because the error deviations without added noises increases with the FFT orders according to Figure 55.
  • As a comparison, for Sin/Cos signal, because Figure 45 shows little increase with the FFT orders of the error deviations without added noises, the improvement of the numerical recovering in Figure 64 is almost independent of FFT orders.
In both cases, error deviations decreases with the strength of the added noise.
9.15 Calculation Response
Figure 65: The FFT execution time (y-axis) versus the FFT order (x-axis) for the implementation using either C++, Java, or Python (legend), with each calculation consisted of $36$ roundtrip FFT transformations. The C++ executable is complied with -O3 -march=native optimization. Without such optimization, the C++ -O0 implementation is about a order of magnitude slower than the Java implementation.
Figure 65 plots the execution time for the roundtrip FFT calculations using either C++, Java, or Python. As expected, in Figure 65, the execution time increases exponentially with the FFT order $L$ as $2^{L}$ for all the three implementations when the FFT order $L>10$, validating the calculation response.
At first, it is surprising that in Figure 65 the Java implementation is 2 times faster than the C++ implementation which has been fully optimized for speed. Because FFT contains a large amount of independent and logically identical branches, the dynamic just-in-time Java compiler [37] can optimize according to data in real-time, compile the bytecodes into native machine codes, and parallelize the execution to the multiple cores and GPUs on a modern computer. Perhaps it is necessary to rethink multi-threading as a way to manually increase the time performance, because the dynamic just-in-time optimization of Java is superior to the static optimization of C++ for repeated and heavy computations with preload data such as FFT.
@@ 9.16 Summary
The Library sine functions implemented using conventional floating-point arithmetic have been shown to contain numerical errors equivalent up to $10^{-3}$ of the input precision in the worst cases of FFT transformations. These $\sin(x)$ numerical errors increase periodically with $x$, potentially resonating with periodic input signals. Computational result obtained with floating-point arithmetic are highly sensitive to the input data, the scale of computation, and the numerical errors in library functions. This implies that a small-scale test cannot reliably qualify or predict the behavior of large-scale calculations, but needs rigorous analysis for the ideal or proper coverage regions in each parameter space. The impact of numerical errors within mathematical libraries has not received sufficient attention. Moreover, such errors in floating-point arithmetic are inherently difficult to detect and quantify.
Variance arithmetic offers a robust alternative, as its computed values closely align with those obtained from conventional floating-point arithmetic while its associated uncertainties systematically trace all input errors. Furthermore, its resulting error deviations provide an objective basis for classifying of calculation quality as ideal, proper, or suspicious.
10 Regressive Generation of Sin and Cos
Figure 67: Uncertainties, value errors (left y-axis), and error deviation (right y-axis) for $\sin(x)^{2}+\cos(x)^{2}-1,x\in[0,\pi/4]$ versus regression order (x-axis) when compared with the Quart sine function (legend)
$\sin(0)=\cos(\frac{\pi}{2})=0;\;\;\;\sin(\frac{\pi}{2})=\cos(0)=1;$ (10.1)
Figure 67 evaluates the Quart and Regression sine functions using $\sin(x)^{2}+\cos(x)^{2}-1=0$. The figure shows that the resulting uncertainties remain nearly constant and independent of the regression order $L$ when $L\geq 4$ for both the Regression sine and the Quart sine functions. It also demonstrates that the value errors for both are comparable. The Regression sine function has error deviations closer to $1$.
11 Conclusion and Discussion
11.1 Summary of Statistical Taylor Expansion
The starting point of statistical Taylor expansion is the assumption that all input variables obeys uncorrelated uncertainties. This is generally regarded as a reasonable and practical statistical requirement for input data [29]. Once this assumption is satisfied, statistical Taylor expansion quantifies uncertainty as the deviation of the value errors and traces track the variable dependencies in intermediate steps using standard statistical methods. While this approach eliminates the dependency problem, it also requires a more rigorous process to determine the resulting mean and deviation for any analytic expression. Furthermore, it can reject invalid calculations on unified statistical and mathematical grounds and explicitly incorporates sample counts and the corresponding underlying uncertainty distributions into the statistical analysis.
11.2 Summary of Variance Arithmetic
Variance arithmetic simplifies statistical Taylor expansion by assuming that the sample counts for all inputs are sufficiently large to achieve ideal variance. Ill-formed problems can invalidate the resulting variance in several ways by producing non-convergent, unstable, negative, infinite, or unreliable results.
The presence of ideal coverage is a necessary condition for a numerical algorithm in variance arithmetic to be considered correct. Ideal coverage also defines the algorithm’s optimal applicable range. Under ideal coverage, the calculated uncertainty equals the deviation of the value errors, and the normalized errors follow either a Normal or Delta distribution depending on the context.
Variance arithmetic provides proper coverage for floating-point rounding errors.
The applicability of variance arithmetic has been demonstrated across a wide range of computational scenarios, including analytic calculations, progressive computation, regressive generalization, polynomial expansion, statistical sampling, and transformations.
The code and analysis framework for variance arithmetic are available as an open-source project at https://github.com/Chengpu0707/VarianceArithmetic.
@@ 11.3 Improvements Needed
This paper presents variance arithmetic which is still in its early stage of development. From a theoretical standpoint, several important questions still need to be addressed.
Library mathematical functions should be recalculated using variance arithmetic, so that each output value is accompanied by its corresponding uncertainty. Without this refinement, the value errors in the library functions can produce unpredictable and potentially significant effects on numerical results. How to compute and validate these new uncertainty-bearing library functions will need innovation. For example, the uncertainty of a 64-bit value can not directly determined by its difference with the corresponding 128-bit value,
Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions. The statistical bounding to find the bounding range $\kappa$ for the bounded moment needs to be extended to discrete distributions.
The choice of ideal bounding range $\hat{\kappa}$ needs further study. For example, the choice of $\hat{\kappa}=5$ for Normal distribution is based on the $5-\sigma$ rule in practice only. Such choice certainly can not be applied to other distributions. Also, the choice of ideal bounding range $\hat{\kappa}$ determines the convergence of both $\log(x)$ and $x^{c}$ as $P(x)\lesssim 1/\hat{\kappa}$, which needs better mathematical understanding. The condition for Formula (2.47) seems to be as generic as the definition in Formula (2.7) which needs further proof.
Cases where $\delta^{2}f<\widehat{\delta^{2}}f$ due to low sample count require further theoretical study and experimental verification.
The dependence of the convergence of $\log(x\pm\delta x)$ and $(x\pm\delta x)^{c}$ on the chose of ideal bounding range $\kappa$ needs further discussion. Determining the uncertainty upper bounds for each analytic expression through analytic methods is an important next step.
The performance of variance arithmetic must be improved for broader practical adoption. The fundamental formulas of statistical Taylor expansion, Formula (2.9), (2.3), (2.12), and (2.3) contain a large number of independent summations, making them excellent candidates for parallel processing. Moreover, the inherently procedural nature of these formulas allows statistical Taylor expansion to be implemented efficiently at the hardware level.
When an analytic expression undergoes statistical Taylor expansion, the resulting expression can become highly complex, as in the case of matrix inversion. However, modern symbolic computation tools such as SymPy can greatly facilitate these calculations. This observation suggests that it may be time to shift from purely numerical programming toward analytic programming, particularly for problems that possess inherently analytic formulations.
@@
A key open question is whether variance arithmetic can be adapted to achieve ideal coverage for floating-point rounding errors. Except for integer multiplication, variance arithmetic does not track rounding errors during computation, because tracking rounding errors for every operation is too expensive to achieve in software. The sensitivity to the library functions increases the difficulty to solve these types of problems. Developing variance arithmetic with ideal coverage for such errors would be quite valuable because many theoretical calculations lack explicit input uncertainties.
Because traditional numerical approaches are based on floating-point arithmetic, they must be reexamined or even reinvented within the framework of variance arithmetic. For instance, most conventional numerical algorithms aim to identify optimal computational paths, whereas variance arithmetic conceptually rejects all path-dependent calculations. Reconciling these two paradigms may present a significant and ongoing challenge.
@@ 12.2 Data Availability Statement
The data set used in this study are all generated in the open-source project at https://github.com/Chengpu0707/VarianceArithmetic. The execution assistance and explanation of the above code are available from the author upon reasonable request.

### v17

@@ 10.9 Ideal Coverage
Table 2 shows the minimal required noise to achieve ideal coverage for FFT transformations at FFT order $L=18$ for different signals and sine functions, which is consistent with the corresponding error distributions in Figure 26, 27, 34, and 35. This shows that an error distribution can qualify its input uncertainty quantification. Without knowing precise result, similar histogram can be constructed from the calculated mean $\overline{f}$ and deviation $\delta f$, and the result data set for $f$. It is worth investigating if such empirical histogram has similar power to reveal tracking of value errors by variance arithmetic.
11 Conclusion and Discussion
@@ 11.1 Summary
When uncorrelated uncertainty condition is met, statistical Taylor expansion outputs the mean, deviation, and reliability of an analytic expression. It tracks the variable dependencies in intermediate steps, rejects invalid calculations, and explicitly incorporates the sample counts and the input uncertainty distributions into the statistical analysis. While statistical Taylor expansion eliminates the dependency problem, it also removes most execution flexibility.
The presence of ideal coverage is a necessary condition for a numerical algorithm using statistical Taylor expansion to be considered correct. The ideal coverage defines the optimal applicable range for an algorithm.
Variance arithmetic simplifies statistical Taylor expansion by introducing numerical rules to elimination invalid results, such as divergent, negative variance, unstable, infinite, or unreliable results. It provides proper coverage for floating-point rounding errors. The applicability of variance arithmetic has been demonstrated across a wide range of computational scenarios.
Library mathematical functions should be recalculated using variance arithmetic, so that each output value is accompanied by its corresponding uncertainty. Without this refinement, the value errors in the library functions can produce unpredictable and potentially significant effects on numerical results.
The code and analysis framework for variance arithmetic are available as an open-source project at https://github.com/Chengpu0707/VarianceArithmetic. A more detailed description of this paper is presented at https://arxiv.org/abs/2410.01223.
@@ 11.2 Improvements Needed
This paper presents statistical Taylor expansion and variance arithmetic, which is still in early stage of development, with several important questions remaining.
Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions. The statistical bounding to find the bounding range $\kappa$ for a sample count needs to be extended to discrete distributions. The choice of ideal bounding range $\hat{\kappa}$ needs to extend to other distributions. The statistical meaning for ideal ratio $\alpha$ in the context of its choice of ideal bounding range $\hat{\kappa}$ needs further investigation: Is $\alpha$ the possibility for the result to be bound by $\hat{\kappa}$?
The performance of variance arithmetic must be improved for broader practical adoption. The fundamental formulas of statistical Taylor expansion, Formula (2.5), (2.3), (2.8), and (2.3), contain a large number of independent summations, making them excellent candidates for parallel processing. Moreover, the inherently procedural nature of these formulas allows statistical Taylor expansion to be implemented efficiently at the hardware level.
A key open question is whether variance arithmetic can be adapted to achieve ideal coverage for floating-point rounding errors, because many theoretical calculations lack explicit input uncertainties. Variance arithmetic does not adjust uncertainty variance when floating-point rounding error occurs during calculation. This leads to error deviations larger than 1. Detecting floating-point rounding error also needs hardware implementation for efficiency.
In variance arithmetic, deviations are comparable with values, but variances are used in calculation. This effectively limits the range of deviations to the square root of that of values. If the sign bit of the floating type can be re-purposed as an exponent bit in a new unsigned floating-point representation, the range of the deviation will be identical to that of values.
When an analytic expression undergoes statistical Taylor expansion, the resulting expression can become highly complex, as in the case of matrix inversion. However, modern symbolic computation tools such as SymPy and Mathematica can greatly facilitate these calculations. This observation suggests that it may be time to shift from purely numerical programming toward analytic programming, particularly for problems that possess inherently analytic formulations.
@@
Because traditional numerical approaches are based on floating-point arithmetic, they must be reexamined or even reinvented within the framework of variance arithmetic. For instance, most conventional numerical algorithms aim to identify optimal computational paths, whereas statistical Taylor expansion conceptually rejects all path-dependent calculations. Reconciling these two paradigms may present a significant and ongoing challenge.
@@
11.3 Possible Connections to Quantum Physics
For input distribution without bounding, Formula (2.3) may not converge, for example, when the input uncertainty is Laplace [4] so that $\zeta(2n)=(2n)!$. Formula (2.3) converges in all cases only when the input uncertainty is bounded due to sampling, which resembles the need for re-normalization in quantum field theory [21]. The quantitative dependence of convergence on the choice of ideal bounding range $\hat{\kappa}$ is similar to the principle of quantum physics: what it is depends on how it is measured. For example, because mathematical divergence seems to correspond to distributional zero, it is even tempting to postulate if a localized experiment changes $\hat{\kappa}$ to produce distributional zero and the quantum result as “waveform collapse” [21]. Both resemblances are worth further investigation.
It is also worthwhile investigating if uncertainty bias produces Lamb shift [21].
@@ 12.2 Data Availability Statement
The data set used in this study are all generated in the open-source project at https://github.com/Chengpu0707/VarianceArithmetic. The execution assistance and explanation of the above code are available from the author upon request.